OpenStudy (anonymous):

John takes 75steps in a minute if the lentgh of each step is 60cms find his speed in km/hr

OpenStudy (dumbcow):

\[(75)(60) \frac{cm}{\min} * \frac{1 km}{100,000 cm} *\frac{60 \min}{1 hr}\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

Really simple. Multiply his speed by the number of steps he takes per minute (75 x 60) then convert the result, which is in cm/min, to km/h